Gucchi:

Solve: the quantity 2x minus 20 divided by 3 = 2x x = −5 x = −2 x = 2 x = 13

Gucchi:

am i right with A?

Gucchi:

@florisalreadytaken

Florisalreadytaken:

i am assuming equation is \( \frac{2x-20}{3}=2x\) as \( 2x\:-\frac{\:20}{3}=\:2x \) would not have a solution you can multiply both sides by 3 to get rid of the fraction: \( \frac{2x-20}{3}\cdot \:3=2x\cdot \:3 \ \ \therefore \ \ 2x-20=6x\) take the \( '+20'\) to th other side: \( 2x=6x+20 \) and then so the same with the variables which inculde `x`: \( 2x-6x=6x+20-6x \implies -4x=20\) finally, divide by \(-4\) and ultimately, you will get \( \boxed{x=-5} \)

Gucchi:

Thank you for this explanation!